Gothan City vehicles have license with 6 characters how many distinct plates are possible if you are allowed to use Capital letters and digits?
Answer #1

number of letters = 26

number of digits = 10

Total = 26+10 = 36

Number of possible 6 character plates = 36*36*36*36*36*36 = 2176782336